# Object-level Geometric Structure Preserving for Natural Image Stitching

### v2

@@ 12 More qualitative results
In this section we provide more qualivative results. We mark images with boxes to indicate misalignment, and use lines (and intersections of lines) to show distortion. Please refer to Fig. 9, 10, 11, 12 and 13. It is shown that our object-level preservation of structures can prevent distortion and misalignment at the same time.
13 Successful cases because of Segment Anything Model
@@
14 Failure cases because of Segment Anything Model
We acknowledge that it’s very important to discuss the situations where, Segment Anything Model [28] fails, leading to the failure of the proposed OBJ-GSP image stitching algorithm. Please refer to Fig. 16 and 17. In the context of our approach, it is not imperative for the SAM [28] to achieve precise segmentation of objects containing semantic information. SAM need only recognize key object contours and boundaries. Therefore, the proposed OBJ-GSP is susceptible to SAM failure only in exceedingly rare instances where features are exceptionally sparse, and objects are highly indistinct. Consequently, SAM’s failure would impact solely the maintenance term of our structure, leading OBJ-GSP to degrade to the performance level of the GSP [7]. Simultaneously, we emphasize that SAM’s failure would result in the nullification of our structural preservation term, causing OBJ-GSP’s performance to regress to that of GSP only under scenarios where SAM proves ineffective. In addressing cases involving distortion and misalignment, we posit that mitigation strategies such as global straightening and multi-bend blending, as employed in Autostitch [3], can be leveraged to alleviate these issues.
@@ 15 Comparasion with UDIS++
UDIS [39](TIP 2021) and UDIS++ [40](ICCV 2023) are a family of attempts to address image stitching problems using deep learning frameworks. We compare with the revised version (UDIS++) as it performs better than UDIS. Like us, UDIS++ also aims to solve distortion issues on top of alignment. In the main text, we compute UDIS++’s NIQE to assess its alignment performance. Since UDIS++ is not feature point based, it cannot calculate MDR, a metric for measuring distortion. Therefore, in the supplementary material, we provide results for four scenes with multiple sets of images to intuitively compare distortion levels. Our method outperforms in both distortion resilience and alignment. Please refer to Fig. 18, 19 and 20.
@@
Figure 11: More qualitative results. Images are from REW [29] dataset. We magnify the ground in the red box and the car in the purple box. OBJ-GSP aligns well but the remaining five methods shows misalignment. Distortion is not observed in this case.

### v4

@@ 12 More qualitative results
In this section we provide more qualivative results. We mark images with boxes to indicate misalignment, and use lines (and intersections of lines) to show distortion. Please refer to Fig. 12, 13, 5, 14 and 15. It is shown that our object-level preservation of structures can prevent distortion and misalignment at the same time.
13 Successful cases because of Segmentation
@@
14 Failure cases
As shown in Fig. 11, OBJ-GSP fails in cases of significant parallax. In the perspective of the left image, the corner of the building appears to be obtuse. However, based on the inference from the right image, the corner of the building should be a right angle. Consequently, the stitching algorithm becomes perplexed, unsure of how to preserve the shape of the house.
In only rare instances OBJ-GSP experiences failures due to the malfunction of SAM [20], such as in conditions characterized by inadequate illumination or an exceedingly sparse set of features. However, we acknowledge that it’s very important to discuss the situations where, Segment Anything Model [20] fails, leading to the failure of the proposed OBJ-GSP image stitching algorithm. Please refer to Fig. 18 and 19. In the context of our approach, it is not imperative for the SAM [20] to achieve precise segmentation of objects containing semantic information. SAM need only recognize key object contours and boundaries. Therefore, the proposed OBJ-GSP is susceptible to SAM failure only in exceedingly rare instances where features are exceptionally sparse, and objects are highly indistinct. Consequently, SAM’s failure would impact solely the maintenance term of our structure, leading OBJ-GSP to degrade to the performance level of the GSP [6]. Simultaneously, we emphasize that SAM’s failure would result in the nullification of our structural preservation term, causing OBJ-GSP’s performance to regress to that of GSP only under scenarios where SAM proves ineffective. In addressing cases involving distortion and misalignment, we posit that mitigation strategies such as global straightening and multi-bend blending, as employed in Autostitch [2], can be leveraged to alleviate these issues.
@@ 15 Comparasion with UDIS++
UDIS [30](TIP 2021) and UDIS++ [31](ICCV 2023) are a family of attempts to address image stitching problems using deep learning frameworks. We compare with the revised version (UDIS++) as it performs better than UDIS. Like us, UDIS++ also aims to solve distortion issues on top of alignment. In the main text, we compute UDIS++’s NIQE to assess its alignment performance. Since UDIS++ is not feature point based, it cannot calculate MDR, a metric for measuring distortion. Therefore, in the supplementary material, we provide results for four scenes with multiple sets of images to intuitively compare distortion levels. Our method outperforms in both distortion resilience and alignment. Please refer to Fig. 20, 21 and 22.